OpenStudy (aaronandyson):

6.4kJ of energy causes a displacement of 64m in a body in the direction of force in 2.5s.Calculate 1.The force applied. 2.Power in HP. Take 1 HP = 746W

OpenStudy (aaronandyson):

@Abhisar @TorilovesVampires

OpenStudy (torilovesvampires):

W = K.E = 6.4 kJ = 6.4 x 103 J W = F.d. cos θ Now the displacement of the body is in the direction of applied force. θ = 0 W = F.d F = W/d = (6.4 x 103 J)/64 m F = 100 N Power P = W/t = (6.4 x 103 J)/(2.5) =16 x 103 W =1.6 x 104 W = (1.6 x 104) / 746 = 0.0021 x 104 H.P. Power P = 21 H.P.

OpenStudy (aaronandyson):

???

OpenStudy (abhisar):

Yo, wassup?

OpenStudy (aaronandyson):

I'm confused. Shouldn't the F be 10.3?

OpenStudy (abhisar):

Work = Force x Displacement and this work done will be equal to energy supplied, Right?

OpenStudy (aaronandyson):

Yes.

OpenStudy (abhisar):

Let the applied force be F, then 6400 = F x 64. Now you can calculate the force.

OpenStudy (aaronandyson):

100N

OpenStudy (abhisar):

You got it ;)

OpenStudy (aaronandyson):

And the P?

OpenStudy (abhisar):

\(\sf Power = \frac{Work/Energy}{Time}\) You know the work, you know the time. Just plug in the values and find power.

OpenStudy (aaronandyson):

64000/25?

OpenStudy (aaronandyson):

2560 W?

OpenStudy (abhisar):

6.4 KJ = 6400 J

OpenStudy (abhisar):

Power = 6400/25

OpenStudy (aaronandyson):

Time is 2.5 s not 25s

OpenStudy (abhisar):

Right, then it's correct!

OpenStudy (abhisar):

Finally, convert the power into HP. 1 HP =746 Watts.

OpenStudy (aaronandyson):

3.43?

OpenStudy (abhisar):

That's right!

OpenStudy (aaronandyson):

Thanks @Abhisar and @TorilovesVampires both.

OpenStudy (abhisar):

pleasure is all mine c: